2. The equation D=1.2√h gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet.
a. Solve this equation for h.
b. Long’s Peak in Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long’s Peak? Can you see Cheyenne, Wyoming (about 89 miles away)? Explain your answer

Answer by nyc_function(2741)

Given D =1.2√h, solve for h.
D/1.2 = √h
[D/1.2]^2 = [√h]^2
D^2/1.44 = h
============================
To answer 2(b), let h = 14,255 and solve equation for D. I'll let you do that.
============================
To answer the second part of 2(b), convert 89 miles to feet (one mile equals 5280 feet) and then replace D with your conversion to solve for h.
I'll let you do that.